# EmbAdvisor: Adaptive Cache Management for Sustainable LLM Serving

### v1

@@ 5.4. Constraint Solver
Notation Description
$N$ Total number of requests in an LLM service
$p_{j}$ Average power at request rate $j$
$T_{j}^{N}$ Total time of serving $N$ requests at request rate $j$
$S_{j}^{N}$ Cache size for serving $N$ requests at request rate $j$
$S_{\rm total}$ Total SSD size
CI Carbon intensity when running LLM service
$\texttt{LT}_{\rm comp}$ Lifetime of $\rm comp\in\{{\rm SSD},{\rm GPU},{\rm CPU},{\rm Mem}\}$
@@
EmbAdvisor minimizes total carbon emission by selecting the optimal cache size while meeting SLOs. The total carbon emissions include both operational carbon and embodied carbon emissions, where embodied emission includes storage (SSD) and non-storage (GPU, CPU, and memory). Since the cache size does not affect the decode phase, we focus on reducing total carbon emissions from the prefill phase. We formulate this as an Integer Linear Programming (ILP) problem. The key notations are summarized in Table 2.
Assuming an LLM service handling $N$ requests with request rate $j$, we note the average power consumption as $p_{j}$ and the serving latency as $T_{j}^{N}$. Given the carbon intensity CI when running the service, operational carbon emission can be calculated by multiplying power, latency and carbon intensity. Because of adaptive caching, SSD embodied carbon emission calculation is different from other hardware components. SSD embodied carbon emission depends on both serving time $T_{j}^{N}$ and dynamic cache size $S_{j}^{N}$. Given the total time of serving $N$ requests at request rate $j$, we calculate the fraction of time these requests occupy SSD’s lifetime. Given cache size, we calculate the utilization of total storage to derive the SSD’s embodied carbon for serving $N$ requests at request rate $j$. Section 5.4 summarizes the objective function and constraints for the ILP problem. For an LLM service with $N$ requests at request rate $j$, EmbAdvisor minimizes total carbon emissions while meeting TTFT and TPOT SLOs.
@@
$\operatorname*{argmin}_{S_{j}^{N}}$ $\bigg{(}\overbrace{p_{j}T_{j}^{N}\texttt{CI}}^{\textbf{Operational carbon}}+\overbrace{\frac{T_{j}^{N}}{\texttt{LT}_{SSD}}\frac{S_{j}^{N}}{S_{\rm total}}C_{\rm e,SSD}}^{\textbf{Cache embodied carbon}}+$
$\overbrace{\sum_{{\rm comp}\in{{\rm GPU},{\rm CPU},{\rm Mem}}}\frac{T_{j}^{N}}{\texttt{LT}_{\rm comp}}C_{\rm e,comp}}^{\textbf{Other embodied carbon}}\bigg{)}$
$s.t.\quad$ $\sum_{i=1}^{N}z_{i}\geq\rho N~{},$
@@
where $z_{i}\!\!\in\!\!\{0,1\}$ is a binary variable such that $z_{i}\!\!=\!\!1$ if request $i$ meets both TTFT and TPOT SLOs, $z_{i}\!\!=\!\!0$ otherwise. $\rho$ specifies the required fraction of prompts that meet the SLOs. We set it as 0.9, i.e., at least 90% prompts should meet SLOs.
We use an off-the-shelf solver (PuLP developers, 2025) to determine the optimal cache configuration, by using Section 5.4 as the objective function. In our experiment, we maintain a number of cache size configurations: 0 (no cache), 1, 2, 4, 8, 12, and 16 TB. EmbAdvisor always takes the optimal cache configuration determined by the ILP solver.
@@ 5.5. Cache Manager and Replacement Policy
The cache controller takes the cache size configuration from the constraint solver and manages the cache. To expand the cache resize, the cache controller allocates more space on the SSDs to save more KV cache of the context. When the cache size shrinks, the cache controller evicts existing cache entries with low scores, until the total cache size reaches the configuration. Then, the spare cache space will be released. Next, we describe the replacement policy, which determines the score of cache entries.
To optimize carbon efficiency, EmbAdvisor introduces a carbon-saving-aware (CSA) cache replacement policy that prioritizes entries based on their potential carbon impact. Unlike conventional cache replacement policies, such as Least Recently Used (LRU), that focus on access recency or frequency, CSA is designed to minimize carbon emissions by evicting cache entries that offer the least carbon savings. As discussed in Section 5.4, carbon emissions from caching depend on both cache size and hit rate. We design a scoring function guided by four key insights:
@@
    Prioritize entries with more hit tokens, which yield higher operational carbon savings, as Section 3.1.1 shows.
@@
    Favor frequently accessed entries, which have higher likelihood of hits and contribute to carbon savings.
@@
    Prefer smaller entries that consume less storage, reducing embodied carbon from SSD usage.
@@
Together, we design the CSA score as follows
(7) $Score=\frac{\#Token\times\#Hit}{Size\times Age}~{}.$
#$\#Token$ is the accumulated hit token number of this cache, representing the volume of reused context. A higher number indicates more operational carbon savings, supporting Insight (i). #$\#Hit$ is the number of cache hits. Frequent access increases the chance of carbon savings by avoiding recomputation, supporting Insight (ii). $Size$ is the cache size of the entry. Dividing by cache size encourages keeping smaller entries that consume less embodied carbon, supporting Insight (iii). $Age$ is how long the cache has stayed in storage, where older entries are less likely to be reused. Penalizing age promotes the eviction of stale entries, supporting Insight (iv).
As described in Section 5.2, we evaluate two LLM tasks. Overall, $Size$ and $Age$ in Equation 7 work in the same way for both tasks, but we specifically adapt #$\#Token$ and #$\#Hit$ fields to each task.
Task 1: Multi-turn conversation. Each turn reuses KV cache of context from previous turns. We adapt Equation 7 to prioritize cache entries that contribute more to carbon savings in this setting:
(8) $Score=\frac{CurTurn\times\#AccuToken}{Size\times Age}~{},$
where $CurTurn$ encourages retaining cache entries deeper in the conversation, since later turns rely more heavily on prior context. #$\#AccuToken$ reflects the total reused tokens across turns, directly tied to operational carbon savings.
@@
(9) $Score=\frac{\#Hit\times AccuDocLen}{Size\times Age}~{},$
where #$\#Hit$ tracks the number of times a document is reused across questions, and $AccuDocLen$ measures the total reused document length, capturing the operational carbon savings.
@@ 6.1. Methodology
We evaluate two models, Llama-3 70B and 8B (Meta, 2024). The 70B model runs on 4×4$\times$4 × NVIDIA L40 GPUs using the platform in Table 1 and the 8B model reduces the GPUs to 2×2$\times$2 × L40 as the model is less demanding. Due to GPU memory constraints, the 70B model is running under INT8. The 8B model runs under the default BF16. Both models have a context window of 8k tokens. When the context goes beyond this limit, we truncate extra context like prior works (Gao et al., 2024; Liu et al., 2024a). We use 4×4$\times$4 × 4 TB Samsung 990 Pro as the SSDs to maintain the maximum cache size of 16 TB. The energy measurement method follows Section 5.2. The caching system is built on top of LMCache (LMCache Team, 2025b), which includes vLLM (Kwon et al., 2023) and continuous batching (Yu et al., 2022) optimizations. We also integrate our carbon-saving-aware CSA replacement policy into LMCache. We use a maximum cache size of 16 TB for the 70B model and a maximum size of 8TB for the 8B model. We evaluate two datasets that correspondent to the two LLM tasks commonly used by prior works (Liu et al., 2024a; Zhong et al., 2024; Gao et al., 2024; Liu et al., 2024b; Chen et al., 2025; Luo et al., 2024; Du et al., 2024; Jeong and Ahn, 2025).

### v2

@@ 5.4. Constraint Solver
$t$ Timestamp during the LLM execution $p_{i}^{j_{t}}$ Average power of request $i$ at request rate $j_{t}$
$\texttt{CI}_{t}$ Carbon intensity (CI) at time $t$ ${\rm TTFT}_{i}^{j_{t}}$ TTFT for request $i$ at request rate $j_{t}$
$j_{t}$ Request rate at time $t$ $S_{i}^{j_{t}}$ Cache size of request $i$ at request rate $j_{t}$
NN Total number of requests in an LLM service $\texttt{LT}_{\rm comp}$ Lifetime of $\rm comp\in\{{\rm SSD},{\rm GPU},{\rm CPU},{\rm Mem}\}$
We first describe the formulation of the optimization problem. Then, we discuss our assumptions and potential error sources that can affect the decisions. Finally, we describe the solver’s implementation.
5.4.1. Problem Formulation
@@
GreenCache minimizes total carbon emissions by selecting the optimal cache size while meeting the SLO attainment goal. The total carbon emissions include both operational carbon and embodied carbon emissions, where embodied emissions include storage (SSD) and non-storage (GPU, CPU, and memory).
The carbon optimization focuses on the prefill phase, where cache hits directly reduce computation by eliminating redundant processing for existing contexts. While cache size does not directly reduce auto-regressive token generation computations in the decode phase, it effectively reduces decode latency by accelerating the prefill phase, mitigating decode delays for the concurrently scheduled requests within a continuous batching system. For this reason, as discussed in Section 2.2, caching influences only the waiting time rather than the total computation of the decode phase.
We formulate this as an Integer Linear Programming (ILP) problem. The key notations are summarized in Table 2. An LLM service handles a total of $N$ requests. At each time $t$, the LLM serving system experiences a request rate of $j_{t}$. Each request $i$ has power consumption $p_{i}^{j_{t}}$, time-to-first-token latency ${\rm TTFT}_{i}^{j_{t}}$, where the request rate of a future time $t$ is predicted using our load predictor. Given the predicted carbon intensity $\texttt{CI}_{t}$ of a future time $t$, the operational carbon emissions for the prefill phase can be calculated by multiplying power, TTFT, and CI. Because of adaptive caching, the calculation of SSD embodied carbon emissions is different from other hardware components, as shown in Equation 4. It depends on both the cache size $S_{i}^{j_{t}}$ and execution time $\rm TTFT_{t}^{j_{t}}$. We calculate the embodied carbon of the SSD capacity allocated to the LLM service.
Then, given the execution time, we calculate the fraction of time a prompt utilizes the SSD’s lifetime. Section 5.4.1 summarizes the objective function and constraints for the ILP problem. For an LLM service with $N$ requests in total with request rate $j_{t}$ at time $t$, GreenCache minimizes total carbon emissions while meeting TTFT and TPOT SLOs as follows:
@@
$\operatorname*{argmin}_{S_{i}^{j_{t}}}$ $\sum_{i=1}^{N}\bigg(\overbrace{p_{i}^{j_{t}}{\rm TTFT}_{i}^{j_{t}}\texttt{CI}_{t}}^{\textbf{Operational carbon}}+\overbrace{\frac{{\rm TTFT}_{i}^{j_{t}}}{\texttt{LT}_{SSD}}S_{i}^{j_{t}}C_{\rm e,SSD}^{\rm Unit}}^{\textbf{Cache embodied carbon}}+\overbrace{\sum_{{\rm comp}\in{{\rm GPU},{\rm CPU},{\rm Mem}}}\frac{t_{\rm TTFT,i}^{j_{t}}}{\texttt{LT}_{\rm comp}}C_{\rm e,comp}}^{\textbf{Other embodied carbon}}\bigg)$
(6) $s.t.\quad$ $\sum_{i=1}^{N}z_{\rm TTFT,i}\geq\rho N\bigwedge\sum_{i=1}^{N}z_{TPOT,i}\geq\rho N~,$
@@
where $z_{\rm TTFT,i},z_{\rm TPOT,i}\!\!\in\!\!\{0,1\}$ are binary variables such that $z_{TTFT,i}\!\!=\!\!1$ if a request $i$ meets the TTFT constraint, and $z_{\rm TPOT,i}\!\!=\!\!1$ means request $i$ meets the TPOT constraint. $\rho$ specifies the required fraction of prompts that meet the requirements. We set it as 0.9, meaning at least 90 % of requests must satisfy both TTFT and TPOT latency requirements, corresponding to the targeted SLO attainment.
5.4.2. Assumptions and Error Analysis
To formulate the optimization problem efficiently, we make three key assumptions. First, we constrain the cache size variable $S_{t}$ to a discrete set of integers, which aligns with the granularity of cache sizes in the cloud. This constraint disallows the selection of fine-grained intermediate values. Second, we assume that the carbon intensity ($\texttt{CI}_{t}$) remains constant within each decision interval (1 hour due to the granularity of the CI dataset (Maji et al., 2023)). Third, we assume that the profiled metrics (e.g., TTFT) reflect the system’s performance under stable conditions. In particular, we collect the offline profiling data after a cache warm-up period using our LCS replacement policy. Therefore, the dependencies between cache sizes, cache hit rate, and the resultant latency are implicitly captured within the profiled results. Satisfying the SLO constraint requires exploring a combinatorial search space, scaling as $O(2^{T})$ in the worst case, where $T$ denotes the count of discrete optimization points defined by the granularity of timestamps. We further prove that this optimization problem is NP-hard via a reduction from the 0–1 Knapsack problem in Appendix A. However, the specific problem in GreenCache is computationally tractable due to the constrained decision space. Unlike the general Knapsack problem, our formulation involves a limited set of discrete cache size candidates (1 TB granularity) and a finite time interval. These practical assumptions effectively prune the search space. We accept the resulting “rounding loss” as a necessary tradeoff to maintain the computational efficiency required for online decision-making.
In addition, we acknowledge three types of errors affecting theoretical optimality. First, the load predictor may introduce errors compared to the ideal outcomes. Second, the CI predictor can also introduce errors. Third, profiling error arises from discrepancies between the profiling phase and actual execution. We evaluate these errors in Section 6.5 and demonstrate that prediction and profiling errors have a low impact on total carbon savings.
5.4.3. Solver Implementation
We use the PuLP optimization modeling library (PuLP developers, 2025) with the COIN-OR CBC solver (COIN-OR Foundation, 2005) as the backend to solve the ILP objective function of Section 5.4.1. GreenCache takes the optimal cache configuration determined by the ILP solver and performs cache resizing every hour. The cache has an allocation granularity of 1 TB, with a maximum size of 16 TB. We evaluate the ILP execution time in Section 6.4 — 7.03 s per decision on average, a low overhead compared to the hourly cache resizing frequency.
@@ 5.5. Cache Manager and Replacement Policy
The cache controller takes the cache size configuration from the constraint solver and manages the cache. To enlarge the cache, the cache controller allocates more SSD space from the cloud. When the cache shrinks, the cache controller evicts existing cache entries with the lowest scores until the total size reaches the configuration. Then, the spare cache space will be released.
The replacement policy determines the score of cache entries. To optimize carbon efficiency, GreenCache introduces Least Carbon Savings (LCS), a new carbon-aware cache replacement policy that prioritizes entries based on their potential carbon impact. Unlike conventional cache replacement policies that focus on either access recency (e.g., LRU) or frequency (e.g., LFU), LCS minimizes carbon emissions by evicting cache entries that offer the least carbon savings. As discussed in Section 5.4, carbon emissions from caching depend on both cache size and hit rate. We design a scoring function guided by four key insights:
@@
    Prioritize entries with more hit tokens, which yield higher operational carbon savings, as shown in Section 3.1.1.
@@
    Favor frequently accessed entries, as they are more likely to produce cache hits and contribute to greater carbon savings.
@@
    Prefer smaller entries that consume less storage, reducing embodied carbon from SSD usage. Although a longer context reduces more operational carbon as concluded in Section 3.1.1, it costs more embodied carbon, which introduces a tradeoff.
@@
Together, we design the LCS score as follows
(7) $\mathit{Score}=\frac{\#\mathit{Token}\times\#\mathit{Hit}}{\mathit{Size}\times\mathit{Age}}~.$
$\mathit{Token}$ is the accumulated hit token number of this cache, representing the volume of reused context. A higher number indicates more operational carbon savings, supporting Insight (i). $\mathit{Hit}$ is the number of cache hits. Frequent access increases the chance of carbon savings by avoiding recomputation, supporting Insight (ii). $\mathit{Size}$ is the cache size of the entry. Dividing by cache size encourages keeping smaller entries that consume less embodied carbon, supporting Insight (iii). $\mathit{Age}$ is how long the cache has stayed in storage, where older entries are less likely to be reused. Penalizing age promotes the eviction of stale entries, supporting Insight (iv).
As described in Section 5.2, we evaluate two LLM tasks. Overall, $\mathit{Size}$ and $\mathit{Age}$ in Equation 7 work in the same way for both tasks, but we specifically adapt $\mathit{Token}$ and $\mathit{Hit}$ fields to each task.
Task 1: Multi-turn conversation. Each turn reuses the KV cache of context from previous turns. We adapt Equation 7 to prioritize cache entries that contribute more to carbon savings:
(8) $\mathit{Score}=\frac{\mathit{CurTurn}\times\#\mathit{AccuToken}}{\mathit{Size}\times\mathit{Age}}~,$
where $\mathit{CurTurn}$ encourages retaining cache entries deeper in the conversation, since later turns rely more heavily on prior context. $\mathit{AccuToken}$ reflects the total reused tokens across turns, directly tied to operational carbon savings.
@@
(9) $\mathit{Score}=\frac{\#\mathit{Hit}\times\mathit{AccuDocLen}}{\mathit{Size}\times\mathit{Age}}~,$
where $\mathit{Hit}$ tracks the number of times a document is reused across questions, and $\mathit{AccuDocLen}$ measures the total reused document length, capturing the operational carbon savings.
@@ 6.1. Methodology
Experiment setup. We evaluate two models, Llama-3 70B and 8B (Meta, 2024). The 70B model runs on $4\times$ NVIDIA L40 GPUs using the platform in Table 1, and the 8B model reduces the GPUs to $2\times$ L40 as the model is less demanding. Due to GPU memory constraints, the 70B model runs under INT8. The 8B model runs under the default BF16. Both models have a context window of 8k tokens. When the context goes beyond this limit, we truncate extra context like prior work (Gao et al., 2024; Liu et al., 2024b). The platform maintains a maximum capacity of 16 TB of SSD. GreenCache provisions SSD at a granularity of 1 TB to the LLM caching system. The power measurement method follows Section 5.2. The caching system is built on top of LMCache (LMCache Team, 2025b), which includes vLLM (Kwon et al., 2023) and continuous batching (Yu et al., 2022) optimizations. We also integrate our carbon-aware LCS replacement policy into LMCache. We use a maximum cache size of 16 TB for the 70B model and 8 TB for the 8B model. By default, GreenCache resizes the cache every hour. Section 6.6.1 evaluates different resizing frequencies.
Tasks and Datasets. We evaluate two datasets that correspond to the two LLM tasks commonly used by prior works (Liu et al., 2024b; Zhong et al., 2024; Gao et al., 2024; Liu et al., 2024a; Chen et al., 2025; Luo et al., 2024; Du et al., 2024; Jeong and Ahn, 2025). In both datasets, we use distinct sets of prompts for profiling and evaluation.
  • Multi-turn conversation based on ShareGPT (ShareGPT, 2023; Hugging Face, 2023). We randomly select a conversation every time and take its next conversation turn as the input prompt. The request follows a Poisson distribution like prior works (Gao et al., 2024, 2025; Zhong et al., 2024). By varying α, the request generator simulates different average request rates. We initialize the cache with 200k prompts.
  • Document comprehension based on TriviaQA (Joshi et al., 2017). Because TriviaQA is initially used for training, it has an almost uniform number of visits per document. Therefore, we introduce skewness by following Zipf distributions similar to prior studies on data caching (Berg et al., 2020; Gill et al., 2007; Hasan et al., 2014; Huang et al., 2013; Breslau et al., 1999). We evaluate Zipf with low and high skewness levels: α=0.4 (10 % of documents are accessed by ~25 % of prompts) and α=0.7 (10 % of documents are accessed by ~50 % of prompts). Prompts are also generated under Poisson distributions. As the context is larger than multi-turn conversations, we initialize the cache with 50k prompts.
Request rate. Prior work has shown that the request rate to LLM services varies in a day (Patel et al., 2024; Stojkovic et al., 2025). To reflect realistic load dynamics, we use the Azure LLM trace (Azure, 2024) to simulate prompt arrivals at corresponding request rates. We downscale the request rate of the Azure trace to match our platform’s capacity, ensuring that the peak rate is still within our system’s sustainable throughput. The load predictor in Section 5.3 forecasts future request rates using past and current rate data.
Carbon intensity. We evaluate four grids, FR, FI, ES, and CISO, and predict their CI using EnsembleCI (Yan et al., 2025), which is trained with 18 months of data and predicts CI for July 6, 2022. The training data and groundtruth CI come from the CarbonCast dataset (Maji et al., 2023). We use the CI on this date for the main performance and carbon evaluation, where we incorporate the variable CI traces with the variable request rate to simulate a realistic scenario. We also use the average CI of the ES grid for sensitivity and ablation studies.
SLOs. We define SLOs for both TTFT and TPOT. For multi-turn conversations, we set TTFT and TPOT SLOs to 2.5 s and 0.2 s for the 70B model, and 0.5 s and 0.15 s for the 8B model. For document comprehension, where inputs are longer and latency is less time-critical, we relax the TTFT thresholds to 15 s for the 70B model and 2.5 s for the 8B model. These SLOs are aligned with prior work (Zhong et al., 2024; Agrawal et al., 2024) and benchmarking results from LLMPerf (LLMPerf, 2024). We focus on an SLO attainment of 90 %, requiring at least 90 % requests to meet the timing constraints.
Comparison points. We evaluate the following system design points: (1) No Cache: Non-caching baseline with vLLM and continuous batching. (2) Full Cache: Use the maximum cache sizes as described in the experiment setup. (3) GreenCache: The carbon-aware caching system in this work. The maximum cache size is the same as Full Cache.
6.2. Carbon Emission and SLO Attainment
Multi-turn conversation.
Document comp. ($\alpha=0.4$).
Document comp. ($\alpha=0.7$).
(a) Carbon emissions of Llama-3 70B.
Multi-turn conversation.
Document comp. ($\alpha=0.4$).
Document comp. ($\alpha=0.7$).
(b) Carbon emissions of Llama-3 8B.
Figure 13. Average carbon emissions of LLM tasks.
Multi-turn conversation.
Document comp. ($\alpha=0.4$).
Document comp. ($\alpha=0.7$).
(a) Llama-3 70B.
Multi-turn conversation.
Document comp. ($\alpha=0.4$).
Document comp. ($\alpha=0.7$).
(b) Llama-3 8B.
Figure 15. SLO attainment timelines of LLM tasks.
We evaluate the carbon emissions and SLO attainment of GreenCache using dynamic CI traces of four grids, where each grid follows an Azure request rate trace.
(a) Multi-turn conversation with Llama-3 70B.
(b) Document comprehension ($\alpha=0.4$) with Llama-3 70B.
Figure 17. Timelines of carbon emissions under variable CI and rate.
Carbon Emissions. First, we present the overall carbon emissions when serving Llama-3 70B and 8B models, as Figures 13(a) and 13(b) show, respectively. Compared to Full Cache, GreenCache achieves average 12.6 %, 9.4 %, and 5.6 % lower carbon in multi-turn conversation and document comprehension with two skewness levels, respectively, when serving Llama-3 70B. In comparison, when serving Llama-3 8B, the carbon emission reductions are slightly lower – 10.8 %, 7.6 %, and 9.7 %, respectively, as the 8B model is lightweight and requires a smaller cache. The carbon emission savings come from a reduced cache usage over Full Cache, as the cache size labels indicate. We also observe that GreenCache achieves higher savings in low-CI grids like FR and FI, 14.4–20.3 %, as embodied carbon is more prominent in these grids. The skewness of the document compression dataset also significantly changes the carbon savings and cache sizes. The high-skewness case ($\alpha=0.7$) requires a smaller cache size compared to the low-skewness case ($\alpha=0.4$) as fewer prompts are frequently used. Thus, the carbon emission savings from GreenCache in the high-skewness case are higher. We notice that although No Cache sometimes achieves lower carbon emissions than GreenCache. However, this option is not acceptable due to the violation of SLOs, which will be discussed next.
SLO Attainment. We evaluate GreenCache’s SLO attainment against baselines in the four grids. Figures 15(a) and 15(b) present the P90 TTFT and TPOT through a day of both models, and compare them against the thresholds as specified by the SLOs. The P90 latency staying below the SLO-specified thresholds indicates at least 90 % SLO attainment. Among all scenarios, GreenCache only exhibits slightly higher P90 latency than Full Cache, staying below both TTFT and TPOT thresholds as specified by the SLOs, indicating over 90 % SLO attainment. In contrast, No Cache’s P90 latency exceeds the SLO constraints frequently, which is not considered a viable solution.
Timelines. Finally, we showcase timelines that demonstrate the dynamics of cache size and per-prompt carbon emissions under real-time CI and request rate for Llama-3 70B. Figures 17(a) and 17(b) show the timelines for the multi-turn conversation and document comprehension (with skewness $\alpha=0.4$) tasks, respectively. The y-axis presents the CI and request rate normalized by their highest values. We do not include the No Cache baseline in the timeline as it fails to meet SLO. Throughout the day, GreenCache reduces carbon emissions by 6.9–20.4 % and 3.2–16.2 % over Full Cache in multi-turn conversation and document comprehension, respectively. Among the 4 grids, FR demonstrates the most carbon reduction due to its low CI, amplifying the impact of reducing embodied carbon; GreenCache reduces carbon emissions by an average of 15.1 % and up to 25.3 % reduction in multi-turn conversation, as GreenCache utilizes 15 TB less SSD than Full Cache (12 AM of the day in Figure 17(a)). Specifically, cache size is more sensitive to the request rate under low CI, as a larger cache is needed to meet the SLO attainment goal. In contrast, under high CI, the cache sizes are generally larger, as caching is effective at reducing operational carbon. For example, in Figure 17(a), FR, FI, and ES grids have similarly small cache sizes, while CISO has more variation in cache sizes. In Figure 17(b), the difference is less prominent as this task has longer contexts and requires larger caches. For the same reason, GreenCache saves less carbon when the load is higher because a larger cache is needed (up to the full 16 TB of cache).
6.3. Ablation Studies
In this section, we first study benefits from the adaptive caching and then compare LCS with other cache replacement policies.
(a) Multi-turn conversation.
(b) Document comp. ($\alpha$=0.4).
(c) Document comp. ($\alpha$=0.7).
Figure 18. Ablation study on adaptive caching (Llama-3 70B).
6.3.1. Adaptive Caching Analysis
We first conduct an experiment that integrates GreenCache’s adaptive caching technique into LMCache (LMCache Team, 2025b) while using its original LRU policy (i.e., LRU + Optimal). Figure 18 shows the carbon reduction under different request rates. For this analysis, we use the average carbon intensity of the ES grid, which is 124 gCO2e/kWh. Compared to Full Cache, GreenCache’s adaptive caching can bring up to 10.3 % carbon savings in multi-turn conversation and 6.6–9.9 % reduction in document comprehension. As the request rate increases, the benefit decreases because GreenCache adapts to larger cache sizes to meet the SLO attainment goal.
6.3.2. Replacement Policy Comparison
We then compare the hit rates of the following replacement policies: (1) First In, First Out (FIFO) replaces blocks in the order in which they were added. (2) Least Recently Used (LRU) replaces the least recently used cache block. It is the default policy in LMCache (LMCache Team, 2025b). (3) Least Carbon Savings (LCS) is the replacement policy in this work that replaces the entry that brings the least carbon emission savings (details in Section 5.5).
We evaluate cache sizes ranging from 1 to 16 TB with the Llama-3 70B model. We define cache hit rate as the number of tokens reused from the cache over the total number of input tokens, as shown in Table 3. In the vast majority of cases, our replacement policy LCS outperforms other replacement policies. Although LRU indicates similar performance as LCS with 16TB cache size, LCS demonstrates up to 9 % higher hit rate than LRU with smaller cache sizes. We also notice that, in the document comprehension task, higher skewness ($\alpha=0.7$) leads to better cache hit rates in all policies, compared to low skewness ($\alpha=0.4$).
Table 3. Hit rate (Llama-3 70B).